OpenStudy (lgbasallote):

besides the gravitational force, a 2.80-kg object is subjected to one other constant force, the object starts from rest and in 1.20 s experiences a displacement of (4.20 i - 3.20 j) m. where the direction of j is the upward vertical direction. Determine the other force.

OpenStudy (lgbasallote):

so what i think i should do...

OpenStudy (lgbasallote):

i think i solve for acceleration first \[a = \frac{2(4.20 \hat i - 3.20 \hat j )}{(1.20)^2}\]

OpenStudy (lgbasallote):

\[a = \frac{2(4.20 i - 3.20 j)}{1.44}\]

OpenStudy (lgbasallote):

so now i use \[\sum y = ma\]

mathslover (mathslover):

K so i think that u need help now : http://answers.yahoo.com/question/index?qid=20080911200716AAYdMDE please have a look if it does not help u then please tell

OpenStudy (lgbasallote):

i mean \[\sum F_y = ma\]

mathslover (mathslover):

@vishweshshrimali5

OpenStudy (lgbasallote):

i dont understand that guy's notations o.O can you explain it to me simply :/

mathslover (mathslover):

ouch ! it hurted me :D may be @vishweshshrimali5 help u

OpenStudy (lgbasallote):

why was that you? lol

OpenStudy (vishweshshrimali5):

Hi @lgbasallote What's the problem

OpenStudy (lgbasallote):

if what im doing is right

OpenStudy (vishweshshrimali5):

How did u calculate the acc. I mean by which formula \[v^2 = u^2 + 2aS\]

OpenStudy (anonymous):

is it a horizontal force?

OpenStudy (vishweshshrimali5):

Is this the formula u used?

OpenStudy (lgbasallote):

i was doing \[F_1 - W = ma\] \[F_1 - (2.8)(9.8) = 2.8 \times \frac{2(4.20 i -3.30j)}{1.44}\]

OpenStudy (lgbasallote):

i used \[a = \frac{2d}{t^2}\]

OpenStudy (anonymous):

is it a horizontal force?

OpenStudy (lgbasallote):

i dont know...

OpenStudy (lgbasallote):

doesnt seem to be

OpenStudy (lgbasallote):

it has i and j

OpenStudy (lgbasallote):

so is my solution right?

OpenStudy (anonymous):

then we will have numerous answers as the force might also act in an angle,considering horizontal force as gravity corresponds to vertical displacement u can find the force acting which corresponds to 4.2m displacemnt in horizontal direction

OpenStudy (lgbasallote):

im guessing the force is wherever the answer is simplest lol..im not a full-time physicist to determine infinite number of forces

OpenStudy (anonymous):

if u simply do s=1/2gt^2 we get 7.2m so that corresponds to 3.2j in the vertical directio so wat does 3.2 correspond to?

OpenStudy (lgbasallote):

1/2gt^2?

OpenStudy (anonymous):

s=ut+1/2gt^2

OpenStudy (lgbasallote):

oh g as i acceleration

OpenStudy (anonymous):

that wud come to 9.45 so s=9.45 and then calculate a then F=ma and hence forcve in HORIZONTAL direction

OpenStudy (lgbasallote):

wait tell me first...was i doing it right

OpenStudy (anonymous):

no for dispalcement u cannot do 4.2-3.2 as they are in different directions

OpenStudy (lgbasallote):

i didnt subtract them o.O

OpenStudy (anonymous):

or u can do like this......resulatnt acceleeration isroot of a^2+g^2

OpenStudy (anonymous):

then resultant displacement is root of 4.2 square +3.2square and then go for finding a

OpenStudy (lgbasallote):

wait....was i doing it right a while ago???/

OpenStudy (anonymous):

u undersatnd why i square and add and take square root as they are mutually perpendicular vectors

OpenStudy (lgbasallote):

i wasnt finished yet anyway

OpenStudy (anonymous):

u get it?

OpenStudy (lgbasallote):

yes i do...what im asking is about my owrk

OpenStudy (lgbasallote):

work*

OpenStudy (anonymous):

no simple subtraction wont fewtch u answer u have to do it vectorially

OpenStudy (lgbasallote):

i wasnt subtracting

OpenStudy (lgbasallote):

it's like x - y i just put it there it's not yet finished

OpenStudy (anonymous):

k go on and finally check ur answer with wat i told

OpenStudy (lgbasallote):

\[F_1 = \frac{2(5.34)}{1.44} + 27.44\] \[F_1 = \frac{10.68}{1.44} + 27.44\] \[F_1 = 7.42 + 27.44\] \[F_1 = 34.86 N\]

OpenStudy (lgbasallote):

is that correct?

OpenStudy (anonymous):

can u expalin ur steps?

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

explain your steps?

OpenStudy (lgbasallote):

i did nothing fancy...it was just algebra..what do you mean explain o.O